A car is travelling at a constant speed of 450m/s. Passes a trooper hidden behind a Billboard one second of the car passes the Billboard, the trooper sets out from the Billboard to catch it, acelerating at a constant speed of 25m/s².how long does it take her to take the car?
1
Expert's answer
2020-02-18T09:40:58-0500

Let the time of motion of the trooper is "t," then distance traveled by her


"d=\frac{at^2}{2}."

The distance traveled by a car


"d=v(t+1)."

Hence

"\frac{at^2}{2}=v(t+1),""\frac{25t^2}{2}=450(t+1),"

"t^2-36t-36=0,"

"t=6 (3 + \sqrt{10})\:\rm s."

Finally, the distance traveled by trooper

"d=\frac{at^2}{2}=\frac{25\times (6 (3 + \sqrt{10}))^2}{2}=462\:\rm m."